A=+68+ a= ?
B= ? b=+8+
C= ? c=+5+
Determine C
A.) +31.38+
B.) +32.92+
C.) +33.12+
D.) +37.02+
Thank you.

Answer by Boreal(15235)

a^2=b^2+c^2-2bc cos A
=64+25-2(40) *0.3746=89-29.97=59.03
sqrt(59.03)=7.68
Now use law sines
sin 68/7.68=sinC/5
sin C=.6034
Take arc sin both sides: C=37.11
I will choose D.